Standard Curve

Test Tube Absorbance [Fe(SCN)2+]
1 0.073 1.0*10-5
2 0.18 2.0*10-5
3 0.212 3.0*10-5
4 0.298 4.0*10-5
5 0.353 5.0*10-5

Finish the table by taking the absorbance of each solution and read the molarity from your standard curve.

Test Tube Absorbance [Fe(SCN)2+]
6 0.132
7 0.238
8 0.341

Please show work!! Thank you!!
Answer #1

Concentration 0.00001 0.00002 0.00003 0.00004 0.00005 Absorbance 0.073 0.18 0.212 0.298 0.353 Chart Title y = 6780x + 0.0198Y = 6780 X +00L18 - equation of line X Concentration y Absorbance, for absonbance 0.132 0.132 -0.0198 = x = [x= 1.6548X106] =6780 For. 0.342 Absorbance 0.341 -0.0198 x = 4.737X18-5 6780 Beern Law = C- Concentration A- Absorbance